At the end of the fiscal year, a company showed the account for temporary investments at $46,750 and the account for inventory as $73,250. For the prior year the temporary investments account was $40,000 and the inventory account was $80,000. Based on this information of increase or decrease that would be shown in the balance sheet with horizontal analysis.
Answer #1

Change in amount ($) :

Temporary investments = 46,750-40,000 = 6,750 increase

Inventory = 73,250-80,000 = -6750 decrease

Percentage change :

Temporary investments = 6750/40000 = 16.88% increase

Inventory = - 6750/73250 = -9.22% decrease